the hypotenuse of a right triangle is 10 cm long. one leg is twice as long as the other. what are the lengths of the legs
Answer by ewatrrr(24785)

 
Hi
Applying the Pythagorean Theorem
x^2 + 2x^2 = 10^2
2x^2 = 100
x^2 = 50 (tossing out negative solution for unit measure)
x = √50, the lengths of shorter leg. Longer 2√50